OpenStudy (anonymous):

Solve the following system using substitution. -3x + y = -2 -x + y = 6

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (mertsj):

You need to fix your first equation.

OpenStudy (anonymous):

haha my bad.

OpenStudy (anonymous):

Saraah, still having troubles?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

-3x + y = -2 -x + y = 6 first, one of the equations needs to be solved for y. Look at the. Which equation is easiest to solve for y?

OpenStudy (compassionate):

...

OpenStudy (anonymous):

the first one

OpenStudy (compassionate):

Well, it doesn't really matter which equation we solve for y, but we need to solve one. So lets solve the second since it's a bit easier. -x + y = 6 y = 6 + x. There! Now we know why y equals. Now all we need to do is SUBSTITUTE it into the other equation. -3x + y = -2 = -3x + 6 + x = -2 Now, all you have to do is solve the equation for x!

OpenStudy (anonymous):

alright

OpenStudy (compassionate):

Did you solve the equation?